OpenStudy (anonymous):

The volume of a sphere with radius r is V(r)=(4/3)pir^3. If a child is adding snow to a perfectly spherical snowball at a rate of 5cm^3/second, at what rate is the radius of the snowball increasing when the radius is 10cm?

OpenStudy (whpalmer4):

Is this a calculus question?

OpenStudy (anonymous):

@whpalmer4 yes, related rates; given \(V=\frac43\pi r^3\) and \(\frac{dV}{dt}=5\); what is \(\frac{dr}{dt}\) when \(r=10\)?

OpenStudy (whpalmer4):

Yeah, just didn't want to suggest a calculus solution if it wasn't a calculus class!

OpenStudy (anonymous):

yeah its calculus

OpenStudy (whpalmer4):

@heradog do you know how to find \(\dfrac{dr}{dV}\)?

OpenStudy (anonymous):

\[dr/dt=dV/dt*dr/dV\] Right? but not sure how to go about it no

OpenStudy (whpalmer4):

Can you solve the volume equation for r?

OpenStudy (anonymous):

4188.79

OpenStudy (whpalmer4):

no, solve for r in terms of V

OpenStudy (anonymous):

This looks like rate of change.

OpenStudy (anonymous):

no or now?

OpenStudy (whpalmer4):

no. we aren't ready for numbers yet.

OpenStudy (anonymous):

this is why I'm so stupid

OpenStudy (whpalmer4):

Nah, you just don't understand it yet, big difference between that and stupid!

OpenStudy (whpalmer4):

We need to find an expression for r in terms of V. When we have that, we will differentiate it with respect to V to give us dr/dV, as a formula. We'll multiply that by our known value of dr/dt and evaluate it at r = 10...

OpenStudy (anonymous):

okay so instead of v= I need r=?

OpenStudy (whpalmer4):

Okay, \[V=\frac{4}{3}\pi r^3\]\[\frac{dV}{dt} = 4\pi r^2\frac{dr}{dt}\](by implicit differentiation) Solve that for \(\frac{dr}{dt}\) and we get \[\frac{dr}{dt} = \frac{dV}{dt}*\frac{1}{4\pi r^2}\] and now we plug in our numbers: \[\frac{dr}{dt} = (5 cm^3/s)*\frac{1}{4\pi (10 cm)^2} = \frac{5}{400\pi}\frac{cm}{s} \approx 0.00398\frac{cm}{s}\]

OpenStudy (whpalmer4):

If we want to check, we can evaluate the volume at the time r = 10 cm: \[V = \frac{4}{3}\pi(10)^3 \approx 4188.79 \]And now after 1 second passes, find the new radius: \[V = \frac{4}{3}\pi r^3\]\[\sqrt[3]\frac{3V}{4\pi} = r\]\[\sqrt[3]{\frac{3*(4188.79+5)}{4\pi}} \approx 10.004\]so our answer looks reasonable.

OpenStudy (anonymous):

Okay cool